OpenStudy (anonymous):

what is 4 5/8 - 2 1/4 - 3 1/2 = ?

OpenStudy (saifoo.khan):

-1 1/8

OpenStudy (anonymous):

the answer sheet says the answer is 5 7/8 but i don't know how to get that answer

OpenStudy (anonymous):

-9/8

OpenStudy (anonymous):

the last sign should be + for that to be the answer

OpenStudy (saifoo.khan):

please chk your ques,there might be a misktae.

OpenStudy (anonymous):

4 5/8 - 2 1/4 + 3 1/2 = ? sorry

OpenStudy (anonymous):

I would rename all of the mixed numbers to improper fractions first. 37/8 - 9/4 -7/2 Then, you will need common denominators. 8 makes sense in this case. 37/8 - 18/8 - 28/8 = -9/8 when you change that to a mixed number, you get -1 1/8

OpenStudy (anonymous):

the sheet answer

OpenStudy (saifoo.khan):

Yeah, its 5 7/8

OpenStudy (anonymous):

37/8 - 18/8 + 28/8 = 47/8 = 5 7/8